OpenStudy (immings):

Iron reacts with dilute HCL to produce hydrogen gas and iron 2 chloride. What volume of dry hydrogen, measured at 22 Celsius and 750 mmHg (room pressure) would be evolved from reacting 0.8827g of iron with an excess amount of HCL?

OpenStudy (aaronq):

You first need to write and balance the reaction for whats happening. Then convert the mass given of Iron to moles. -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. -------------------------------------------------- Since the other reactant is in excess, the limiting reactant is Iron. So use the moles of iron for find the moles of \(H_{2~ (g)}\) produced using the stoichiometric coefficients. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- Finally, use moles found of hydrogen gas and the ideal gas formula: PV=nRT Solve for volume.